OpenStudy (anonymous):

Please I am desperate! Check my answer. Question: Which equation would make this system consistent and dependent? Equation: 2/3y = 7x - (1/2). Choices: A. 42x - 4y = 3 B. 21x = 2y - 3 D. y = (21/2)x - (3/2) My Answer: D

OpenStudy (jtvatsim):

A system is consistent and dependent if the two equations are identical. So we must find an answer choice that is the same as the original equation.

OpenStudy (jtvatsim):

As it is written, our original equation is very messy, using some algebra, we can write: (2/3)y = 7x - (1/2) as 2y = 21x - (3/2) y = (21/2)x - (3/4) Answer choice D is close, but not quite right (it has 3/2 instead of 3/4).

OpenStudy (jtvatsim):

If we solve answer B for y: 21x + 3 = 2y (21/2)x + (3/2) = y which is not the same as the original.

OpenStudy (jtvatsim):

That leaves us with only one answer left: A. 42x - 4y = 3 42x = 3 + 4y 42x - 3 = 4y (42/4)x - (3/4) = y reversing the equation to make the y look right (we can do this since it is equal) y = (42/4)x - (3/4) y = (21/2)x - (3/4) Which is the same as our original equation and we are done!

OpenStudy (anonymous):

Wow, thank you, that makes so much more sense.

OpenStudy (jtvatsim):

no problem, good luck!